# GenAI vs. Human Creators: Procurement Mechanism Design in Two-/Three-Layer Markets

### v1

@@ 2.4.1 Revenue-Maximizing Procurement Mechanism $\mathcal{M}_{3}$ with a Creator Union
Pessimistically, general high-dimensional mechanism design problems remain open and agnostic (briest2010pricing; hart2013menu; daskalakis2015multi; hart2017approximate). Therefore, we formulate the corresponding optimization problem and derive bounds on platform revenue and social welfare, leaving explicit analytical solutions for future research. The revenue-maximizing platform now faces the following optimization problem (Mechanism $\mathcal{M}_{3}$), where the payment rule $p(\cdot)$ is a scalar function and the corresponding social welfare is $SW_{3}={\mathbb{E}}_{c}[-\sum_{k=1}^{K}c_{k}x_{k}^{\rho_{k}}(c)]$:
@@
The main challenge in this problem is that the union may misreport multiple components of the cost vector $c$ simultaneously. The resulting correlations make the optimization problem effectively contain infinitely many constraints and require the use of path integrals in defining the payment rule. Finding the optimal solution remains a difficult open problem in economics. In contrast, as shown in the next section, a platform that maximizes social welfare can efficiently obtain the optimal solution, highlighting a notable advantage.
@@ 2.4.2 Welfare-Maximizing Procurement Mechanism $\mathcal{M}_{4}$ under a Creator Union
 Calculate the payment: $p(\widehat{c})=\widehat{c}\cdot y+\int_{\widehat{c}}^{b}y(t)\cdot dt$ [^5] The symbol $\cdot$ represents the dot product..
@@
This theorem reveals a surprising result: despite the high-dimensional setting, the welfare-maximizing mechanism can be computed efficiently. The proof is constructive, identifying an optimal payment rule $p(\cdot)$ that is independent of the integration path. The distinctive properties of GenAI are crucial to this result. Since GenAI introduces only convex demand constraints, the allocation, whether at a vertex or along a smooth segment, remains orthogonal to cost after transformation. We construct a potential function $\Psi(c)$, detailed in Section˜C.5, and establish the relationship between its gradient and the allocation rule. This approach provides a new method for addressing high-dimensional mechanism design problems, distinct from cycle monotonicity (lavi2007truthful), and advances understanding of this open question. Building on the above results, the next section presents a partial comparison of $Rev_{1-4}$ and $SW_{1-4}$.
@@ 2.5 Comparison of Revenue and Social Welfare for Mechanisms $\mathcal{M}_{1-4}$
The most technically interesting comparison lies between $Rev_{2}$ and $Rev_{4}$. Since the payment rule of $\mathcal{M}_{4}$ involves a path integral, we select a specific integration path and, using the validity of the mechanism, show that $Rev_{2}\geq Rev_{4}$. The complete proof is provided in Section˜C.6.
@@ 3 Content Procurement in Platform-Broker-Creator Markets
In this section, we extend our analysis beyond the union case and consider a three-layer market (fallah2024three). In this setting, the platform first offers a contract $(z,t)$ to a data broker (liu2021dealer), who decides whether to accept it. The broker then engages with human creators and faces a mechanism design problem similar to that in Section˜2.1. Unlike the union, which maximizes the total utility of creators, the broker focuses solely on profit, defined as the difference between $z$ and $\sum_{k=1}^{K}p_{k}$. In practice, companies such as SchoolDigger and Datarade serve as examples of data brokers (zhang2024survey). To build intuition, Figure˜7 in Section˜D.1 illustrates the timeline of this mechanism design setting.
@@ 3.1 Revenue-Maximizing Procurement Mechanism $\mathcal{M}_{5}$ with a Data Broker
We first consider the scenario in which the platform seeks to maximize its revenue. To identify the optimal mechanism, we begin by formulating the broker’s optimization problem given $(z,t)$. Since both the allocation and payment rules depend on $(z,t)$, we denote them by $x(\cdot;z,t)$ and $p(\cdot;z,t)$, respectively. Because the broker signs the contract with the platform in advance, the credibility constraint (https://doi.org/10.3982/ECTA15925) imposes an additional condition that, for any report $\widehat{c}$,
@@ 3.3 The Lose-Lose Effect in Three-Layer Markets
Finally, we compare the platform’s revenue and social welfare in the three-layer market with those in the two-layer setting. We show that both revenue and social welfare decline relative to $\mathcal{M}_{2}$ and $\mathcal{M}_{4}$, indicating that the presence of data brokers distorts allocations and reduces market efficiency. Public contracting further hinders the effective incentivization of downstream creators and prevents the platform from optimizing social welfare based on realized types. This finding highlights the growing need for government oversight and regulation of data markets (AIWebBreak2024).
@@ 4 Empirical Studies
Figure 4: Concrete revenue and social welfare of different mechanisms.
@@
Finally, we simulate the difference between scenarios with and without GenAI, that is, with $\mu=0$. Using the same setup as the preceding evaluation and mechanism $\mathcal{M}_{1}$ as an illustrative example, we report detailed results in Table˜2 of Appendix˜E. In our framework, GenAI leverages human-created content to generate additional material, allowing platforms to meet total demand with less direct human creation. Consequently, both revenue and social welfare increase in the presence of GenAI. Specifically, the sample-averaged revenue rises from –30.90 to –16.52, and sample average social welfare improves from –5.66 to –4.47, representing gains of 46.54% and 21.02%, respectively. These results indicate that GenAI enhances production efficiency while increasing both platform revenue and social welfare. However, as the primary beneficiary of GenAI, the platform enjoys a substantially larger increase in revenue than the corresponding improvement in social welfare. Although the magnitude of improvement depends on parameter initialization, the asymmetric growth between revenue and welfare reflects a broader pattern observed in practice (hosseini2025distributive), highlighting reflections regarding the distributional impact of GenAI adoption.
@@ 5 Conclusion and Discussion
In this paper, we study non-linear procurement mechanism design in markets where human creators coexist with GenAI. We develop optimal mechanisms for platforms seeking to maximize either revenue or social welfare across three market structures: two-layer markets (with or without a creator union) and three-layer markets involving data brokers. Our analysis shows that although the rapid advancement of GenAI substantially affects creators in certain domains, it cannot fully replace them in competitive environments (chiang2023can). Finally, we uncover a striking result: the three-layer market structure leads to a lose-lose outcome, echoing growing societal concerns over AI regulation and governance (smuha2021race; de2021artificial; minssen2023challenges).
Questions naturally arise for future exploration. For general high-dimensional mechanism design problems, such as Mechanism $\mathcal{M}_{3}$, can the optimal solution be computed exactly in polynomial time? Given that our optimization problem retains partial linearity, could approximate solvers be effective when parameters $\rho$ and $\gamma$ are unknown? Since the three-layer market leads to simultaneous declines in both revenue and social welfare, how should government intervention be structured, for instance, what constitutes an appropriate tax rate for data brokers? Furthermore, does competition among human creators yield excess profits for data intermediaries? In light of the growing demand for data in the AI era and the distinctive nature of data as an economic good, these questions open several promising directions for future exploration.
@@ A.1 Related Works
The topic of optimal mechanism design was introduced by the seminal paper myerson1981optimal. Over decades of academic attention, researchers have not only focused on optimal mechanism design from the seller’s perspective but have also begun to consider it from the buyer’s perspective, especially when buyers are either significantly large or even monopolistic, leading to studies on optimal procurement mechanism design (mcafee1987auctions; laffont1993theory). In the economics literature, burguet2012limited examines mechanism design in procurement under limited liability and explores ways to reduce the sponsor’s cost, while gerding2010scalable addresses the service procurement problem in the presence of uncertain duration. garg2008mechanism adopts the framework of a Stackelberg game to study procurement auction design. In management science and supply chain literature, zhang2010procurement investigates procurement with price-sensitive demand, while prasad2013mechanism; ketankumar2015green focus on the procurement of cloud computing resources. jin2002procurement; huang2011mechanism attempt to design optimal procurement auctions to enhance suppliers’ profit margins. Our paper goes beyond traditional procurement mechanism design for physical goods by focusing on the digital content market, which features transferability. This extends the boundaries of procurement mechanism design in this domain.
@@
The development of generative AI has brought AI-generated content into the economic market, sparking research on the nature of AI products and traditional goods. liu2024llm discusses the impact of large language models (LLM) on traditional search engines, while zhang2024llm explores advertising business in the context of LLMs. iyer202413 conducts a case study on the role of LLMs in e-commerce. chkirbene2024large provides a comprehensive survey of GenAI’s applications, challenges, and trends from several aspects, including content creation and marketing. Meanwhile, due to the enormous data demand for GenAI training, public datasets are becoming insufficient, leading companies to shift toward private datasets, which has raised concerns about data pricing. Data Shapley and its variants are the most commonly used methods for data pricing (jia2019towards; ghorbani2019data; schoch2022cs), while recent ai2024instrumental proposes a novel data pricing method based on instrumental value. Our paper studies how much payment is needed to incentivize human creators to generate sufficient content within an agent-based framework, broadening the perspective on data pricing related to GenAI.
@@
With the rapid development of generative AI, competition between human creators and GenAI has gradually emerged (esmaeili2024strategize). In contrast to traditional automation (acemoglu2019automation), GenAI relies on human-created data for training, thereby simultaneously complementing and competing with human labor. yao2024human employs the framework of a Tullock contest to examine the symbiosis and conflict between humans and generative AI. immorlica2024generative examines the augmentation of users by AI agents and the corresponding shifts in equilibrium. gao2025pandora, within a game-theoretic framework, investigates the impact of AI tools on market outcomes and consumer welfare. It highlights that in certain scenarios, both platforms and creators may be reluctant to adopt GenAI. Moreover, some papers discuss the unique economic characteristics of LLMs, or broadly GenAI; please refer to fish2024algorithmic; duetting2024mechanism for further details. Meanwhile, some papers explore the collaboration between humans and GenAI in various tasks (fui2023generative; singh2023human; han2024teams). We adopt a mechanism design perspective to study the collaboration and competition between human creators and GenAI. GenAI relies on human-created content for training, while the content generated by AI partially substitutes human-created content, providing a novel approach to studying the competition between humans and AI.
@@
The three-layer market offers a more comprehensive framework for characterizing data markets compared to the two-layer market. For further details, zhang2024survey provides a survey on data markets. fallah2024three models the three-layer data market as a multistage game and focuses on the subgame Nash equilibrium, while xu2020data considers a three-layer Stackelberg game in the car-sharing market. Besides, reinders2018exploration; song2019contract consider three-layer structures in the domain of supply chain. wang2023privacy investigates privacy protection issues in a blockchain-based P2P three-layer market. balseiro2021multistage; zeithammer2024auctions investigate the impact of intermediaries on the payoffs of buyers and sellers in auctions, as well as the phenomenon of double-shading in the market. We compare the impacts of two-layer and three-layer markets, within the framework of competition between humans and GenAI, on both platform revenue and social welfare. We theoretically demonstrate that the three-layer structure would lead to a lose-lose outcome, offering a novel perspective for future AI data regulation.

### v2

@@ 2.4.1 Revenue-Maximizing Procurement Mechanism $\mathcal{M}_{3}$ with a Creator Union
Pessimistically, general high-dimensional mechanism design problems remain open and agnostic (Briest et al., 2010; Hart et al., 2013; Daskalakis, 2015; Hart and Nisan, 2017). Therefore, we formulate the corresponding optimization problem and derive bounds on platform revenue and social welfare, leaving explicit analytical solutions for future research. The revenue-maximizing platform now faces the following optimization problem (Mechanism $\mathcal{M}_{3}$), where the payment rule $p(\cdot)$ is a scalar function and the corresponding social welfare is $SW_{3}={\mathbb{E}}_{c}[-\sum_{k=1}^{K}c_{k}x_{k}^{\rho_{k}}(c)]$:
@@
The main challenge in this problem is that the union may misreport multiple components of the cost vector $c$ simultaneously. The resulting correlations make the optimization problem effectively contain infinitely many constraints and require the use of path integrals in defining the payment rule. Finding the optimal solution remains a difficult open problem in economics. To be specific, recall in the one-dimensional case, the optimal payment in Algorithm˜1 has the form $p_{k}=c_{k}y_{k}+\int_{c}^{b}y_{k}(c_{1},...,c_{k-1},t_{k},c_{k+1},...,c_{K})dt_{k}$ with truthful reports. However, in the high-dimensional case, the corresponding payment would become $p=c\cdot y+\int_{c}^{b}y(t)\cdot t$ [^5] The symbol $\cdot$ represents the dot product.. When $c$ is a high-dimensional vector, it raises an extra question that the integration results might be different for different paths from $c$ to $b$, which contradicts the fixed contract. Hence, simply applying methods in the one-dimensional case won’t yield a valid mechanism, and it hinders the optimal high-dimensional mechanism design in relevant scenarios. In contrast, as shown in the next section, a platform that maximizes social welfare can efficiently obtain the optimal solution, highlighting a notable advantage.
@@ 2.4.2 Welfare-Maximizing Procurement Mechanism $\mathcal{M}_{4}$ under a Creator Union
 Calculate the payment: $p(\widehat{c})=\widehat{c}\cdot y+\int_{\widehat{c}}^{b}y(t)\cdot dt$.
@@
This theorem reveals a surprising result: despite the high-dimensional setting, the welfare-maximizing mechanism can be computed efficiently. The proof is constructive, identifying an optimal payment rule $p(\cdot)$ that is independent of the integration path. The distinctive properties of GenAI are crucial to this result. Since GenAI introduces only convex demand constraints, the allocation, whether at a vertex or along a smooth segment, remains orthogonal to cost after transformation. We construct a potential function $\Psi(c)$, detailed in Section˜C.6, and establish the relationship between its gradient and the allocation rule. This approach provides a new method for addressing high-dimensional mechanism design problems, distinct from cycle monotonicity (Lavi and Swamy, 2007), and advances understanding of this open question. Building on the above results, the next section presents a partial comparison of $Rev_{1-4}$ and $SW_{1-4}$.
@@ 2.5 Comparison of Revenue and Social Welfare for Mechanisms $\mathcal{M}_{1-4}$
The most technically interesting comparison lies between $Rev_{2}$ and $Rev_{4}$. Since the payment rule of $\mathcal{M}_{4}$ involves a path integral, we select a specific integration path and, using the validity of the mechanism, show that $Rev_{2}\geq Rev_{4}$. The complete proof is provided in Section˜C.7.
@@ 3 Content Procurement in Platform-Broker-Creator Markets
In this section, we extend our analysis beyond the union case and consider a three-layer market (Fallah et al., 2024). In this setting, the platform first offers a contract $(z,t)$ to a data broker (Liu et al., 2021), who decides whether to accept it. The broker then engages with human creators and faces a mechanism design problem similar to that in Section˜2.1. Unlike the union, which maximizes the total utility of creators, the broker focuses solely on profit, defined as the difference between $z$ and $\sum_{k=1}^{K}p_{k}$. In practice, companies such as SchoolDigger and Datarade serve as examples of data brokers (Zhang et al., 2024a). To build intuition, Figure˜7 in Section˜D.1 illustrates the timeline of this mechanism design setting.
@@ 3.1 Revenue-Maximizing Procurement Mechanism $\mathcal{M}_{5}$ with a Data Broker
We first consider the scenario in which the platform seeks to maximize its revenue. To identify the optimal mechanism, we begin by formulating the broker’s optimization problem given $(z,t)$. Since both the allocation and payment rules depend on $(z,t)$, we denote them by $x(\cdot;z,t)$ and $p(\cdot;z,t)$, respectively. Because the broker signs the contract with the platform in advance, the credibility constraint (Akbarpour and Li, 2020) imposes an additional condition that, for any report $\widehat{c}$,
@@ 3.3 The Lose-Lose Effect in Three-Layer Markets
Lastly, we compare the platform’s revenue and social welfare in the three-layer market with those in the two-layer setting. We show that both revenue and social welfare decline relative to $\mathcal{M}_{2}$ and $\mathcal{M}_{4}$, indicating that the presence of data brokers distorts allocations and reduces market efficiency. Public contracting further hinders the effective incentivization of downstream creators and prevents the platform from optimizing social welfare based on realized types. This finding highlights the growing need for government oversight and regulation of data markets (Brooks, 2024).
@@
We end this section by noting that although the proof of the lose-lose effect relies on the specific model, we conjecture that it is a pervasive phenomenon in three-layer markets. The fundamental reason for the lose-lose effect is as follows. When a platform and a broker sign a public contract for content procurement, such as the agreement between OpenAI and The Wall Street Journal, such large-scale deals are typically difficult to keep private due to regulatory requirements. As a result, creators in the relevant domain anticipate higher demand for their content and raise their payment requirements. This weakens the platform’s ability to price discriminate among human creators and naturally reduces the platform’s revenue. For social welfare, once human creators become aware of the demand specified in the broker’s contract, the broker must not only compensate them for their production costs but also incentivize them to generate at least a target level of content. This requires the broker to provide additional informational rents to ensure participation and effort. Since social welfare depends solely on the true production costs, these additional incentives introduce distortions in the allocation. As a result, the maximization of social welfare is distorted, leading to lower social welfare in the three-layer market compared to the two-layer market.
@@ 4 Empirical Studies
Figure 4: Concrete revenue and social welfare of different mechanisms (red: mean; orange: median).
@@
Finally, we simulate the difference between scenarios with and without GenAI, that is, with $\mu=0$. Using the same setup as the preceding evaluation and mechanism $\mathcal{M}_{1}$ as an illustrative example, we report detailed results in Table˜2 of Appendix˜E. In our framework, GenAI leverages human-created content to generate additional material, allowing platforms to meet total demand with less direct human creation. Consequently, both revenue and social welfare increase in the presence of GenAI. Specifically, the sample-averaged revenue rises from –30.90 to –16.52, and sample average social welfare improves from –5.66 to –4.47, representing gains of 46.54% and 21.02%, respectively. These results indicate that GenAI enhances production efficiency while increasing both platform revenue and social welfare. However, as the primary beneficiary of GenAI, the platform enjoys a substantially larger increase in revenue than the corresponding improvement in social welfare. Although the magnitude of improvement depends on parameter initialization, the asymmetric growth between revenue and welfare reflects a broader pattern observed in practice (Hosseini and Khanna, 2025), highlighting reflections regarding the distributional impact of GenAI adoption.
@@ 5 Conclusion and Discussion
In this paper, we study non-linear procurement mechanism design in markets where human creators coexist with GenAI. We develop optimal mechanisms for platforms seeking to maximize either revenue or social welfare across three market structures: two-layer markets (with or without a creator union) and three-layer markets involving data brokers. Our analysis shows that although the rapid advancement of GenAI substantially affects creators in certain domains, it cannot fully replace them in competitive environments (Chiang and Lee, 2023). Finally, we uncover a striking result: the three-layer market structure leads to a lose-lose outcome, echoing growing societal concerns over AI regulation and governance (Smuha, 2021; de Almeida et al., 2021; Minssen et al., 2023).
This work leaves open many interesting future directions. For general high-dimensional mechanism design problems, such as Mechanism $\mathcal{M}_{3}$, can the optimal solution be computed exactly in polynomial time? Given that our optimization problem retains partial linearity, could approximate solvers be effective when parameters $\rho$ and $\gamma$ are unknown? Since the three-layer market leads to simultaneous declines in both revenue and social welfare, how should government intervention be structured, for instance, what constitutes an appropriate tax rate for data brokers? Furthermore, does competition among human creators yield excess profits for data intermediaries? Finally, our static model leads to predictions and insights about the ultimate equilibrium outcomes. How will these insights change in dynamic settings when GenAI and human creators dynamically interact? In light of the growing demand for data in the AI era and the distinctive nature of data as an economic good, these questions open several promising directions for future exploration.
@@ A.1 Related Works
The topic of optimal mechanism design was introduced by the seminal paper Myerson (1981). Over decades of academic attention, researchers have not only focused on optimal mechanism design from the seller’s perspective but have also begun to consider it from the buyer’s perspective, especially when buyers are either significantly large or even monopolistic, leading to studies on optimal procurement mechanism design (McAfee and McMillan, 1987; Laffont and Tirole, 1993). In the economics literature, Burguet et al. (2012) examines mechanism design in procurement under limited liability and explores ways to reduce the sponsor’s cost, while Gerding et al. (2010) addresses the service procurement problem in the presence of uncertain duration. Garg and Narahari (2008) adopts the framework of a Stackelberg game to study procurement auction design. In management science and supply chain literature, Zhang (2010) investigates procurement with price-sensitive demand, while Prasad and Rao (2013); Ketankumar et al. (2015) focus on the procurement of cloud computing resources. Jin and Wu (2002); Huang et al. (2011) attempt to design optimal procurement auctions to enhance suppliers’ profit margins. Our paper goes beyond traditional procurement mechanism design for physical goods by focusing on the digital content market, which features transferability. This extends the boundaries of procurement mechanism design in this domain.
@@
The development of generative AI has brought AI-generated content into the economic market, sparking research on the nature of AI products and traditional goods. Liu et al. (2024a) discusses the impact of large language models (LLM) on traditional search engines, while Zhang et al. (2024b) explores advertising business in the context of LLMs. Iyer et al. (2024) conducts a case study on the role of LLMs in e-commerce. Chkirbene et al. (2024) provides a comprehensive survey of GenAI’s applications, challenges, and trends from several aspects, including content creation and marketing. Meanwhile, due to the enormous data demand for GenAI training, public datasets are becoming insufficient, leading companies to shift toward private datasets, which has raised concerns about data pricing. Data Shapley and its variants are the most commonly used methods for data pricing (Jia et al., 2019; Ghorbani and Zou, 2019; Schoch et al., 2022), while recent Ai et al. (2024) proposes a novel data pricing method based on instrumental value. Our paper studies how much payment is needed to incentivize human creators to generate sufficient content within an agent-based framework, broadening the perspective on data pricing related to GenAI.
@@
With the rapid development of generative AI, competition between human creators and GenAI has gradually emerged (Esmaeili et al., 2024). In contrast to traditional automation (Acemoglu and Restrepo, 2019), GenAI relies on human-created data for training, thereby simultaneously complementing and competing with human labor. Yao et al. (2024) employs the framework of a Tullock contest to examine the symbiosis and conflict between humans and generative AI. Immorlica et al. (2024) examines the augmentation of users by AI agents and the corresponding shifts in equilibrium. Gao et al. (2025), within a game-theoretic framework, investigates the impact of AI tools on market outcomes and consumer welfare. It highlights that in certain scenarios, both platforms and creators may be reluctant to adopt GenAI. Moreover, some papers discuss the unique economic characteristics of LLMs, or broadly GenAI; please refer to Fish et al. (2024); Duetting et al. (2024) for further details. Meanwhile, some papers explore the collaboration between humans and GenAI in various tasks (Fui-Hoon Nah et al., 2023; Singh et al., 2023; Han et al., 2024). We adopt a mechanism design perspective to study the collaboration and competition between human creators and GenAI. GenAI relies on human-created content for training, while the content generated by AI partially substitutes human-created content, providing a novel approach to studying the competition between humans and AI.
@@
The three-layer market offers a more comprehensive framework for characterizing data markets compared to the two-layer market. For further details, Zhang et al. (2024a) provides a survey on data markets. Fallah et al. (2024) models the three-layer data market as a multistage game and focuses on the subgame Nash equilibrium, while Xu et al. (2020) considers a three-layer Stackelberg game in the car-sharing market. Besides, Reinders et al. (2018); Song and He (2019) consider three-layer structures in the domain of supply chain. Wang et al. (2023) investigates privacy protection issues in a blockchain-based P2P three-layer market. Balseiro et al. (2021); Zeithammer and Choi (2024) investigate the impact of intermediaries on the payoffs of buyers and sellers in auctions, as well as the phenomenon of double-shading in the market. We compare the impacts of two-layer and three-layer markets, within the framework of competition between humans and GenAI, on both platform revenue and social welfare. We theoretically demonstrate that the three-layer structure would lead to a lose-lose outcome, offering a novel perspective for future AI data regulation.
A.2 Relative Terms
We use this subsection to illustrate some terminology used in this paper for readers with less economic background. These terms are definitional, but we also provide the economic intuition behind them to assist the reader’s understanding.
  • $\bullet$
    Virtual costs: A variant of the original cost, formally defined as $c+\frac{F}{f}$. Because the platform must incentivize creators to report their true costs, the effective cost the platform faces is the human creator’s production cost $c$ plus the incentive-related virtual cost $\frac{F}{f}$.
  • $\bullet$
    Reverse hazard rates: Originating from survival analysis (Kalbfleisch and Prentice, 2002), and defined in this paper as $\frac{f}{F}$. It is used to characterize the log-concavity of $F$.
  • $\bullet$
    Incentive compatibility: The terminology for a desirable mechanism property that for any creator with private cost $c$, truthfully reporting $c$ to the platform always maximizes the creator’s expected utility (payment minus cost). Hence incentive incentive-compatible mechanisms can help avoid dishonest agent behaviors during information elicitation.
  • $\bullet$
    Individual rationality: The terminology of another desirable mechanism property that human creators always have non-negative expected utility (so they prefer to stay in the market than to exit).
  • $\bullet$
    Path independence: When payments are defined through a path integral (e.g., $\int_{\widehat{c}}^{b}y(t)\cdot dt$ in Mechanism $\mathcal{M}_{4}$), we require that the integral yields the same value for any path from $\widehat{c}$ to $b$.
  • $\bullet$
    Myerson’s analysis: This refers to a classic and well-known analysis framework of Myerson (1981) that converts payments into integrals of the allocation functions, which reduces mechanism design to allocation rule design.
  • $\bullet$
    Convex procurement: We hope the feasible set induced by demand constraints to be convex, so that the optimal mechanism can be computed efficiently in practice.
  • $\bullet$
    Valid mechanisms: A creator’s higher cost should lead the platform to procure less from them, while other creators’ higher costs should lead the platform to procure more from this creator. This reflects substitutability in production.